If we build a barometer using hexane (density equal to 0.655 g/mL) calculate the height of the equivalent column that exerts the same pressure as 76 cm of Hg
1
Expert's answer
2020-12-03T13:53:35-0500

pressure = dHg*g*hHg = dHexane*g*hHexane (d is density); dHg = 13.6 g/mL

dHg*hHg = dHexane*hHexane

hHexane=dHg*hHg/dHexane = (13.6 g/mL)*(76 cm)/(0.655 g/mL) = 1578 cm